# Offline Learning for Combinatorial Multi-armed Bandits

### v1

@@ Appendix A Proof for the Upper Bound Result
Appendix B Proof for the Lower Bound Result
@@
Recall that $A(\mathcal{D})$ is the action returned by algorithm $A$ and we use the Le Cam’s method [122]:
@@
where inequality (a) is due to Lemma 8 and inequality (b) comes from: $\mathrm{KL}\left(\mathbb{P}_{\bm{p}\otimes\bm{\mu}_{1}}\|\mathbb{P}_{\bm{p}\otimes\bm{\mu}_{2}}\right)\leq\frac{n\mathrm{KL}\left(\mathbb{P}_{\bm{\mu}_{1}}\|\mathbb{P}_{\bm{\mu}_{2}}\right)}{C_{\infty}^{\star}}\leq\frac{n(2\Delta)^{2}}{C_{\infty}^{\star}\left(1/4-\Delta^{2}\right)}\leq 20n\Delta^{2}/C_{\infty}^{\star}$. Here we use the fact that each arm in the path are fully dependent Bernoulli random variables, $\mathrm{KL}\left(\mathrm{Bern}(p)||\mathrm{Bern}(q)\right)\leq\frac{(p-q)^{2}}{q(1-q)}$ and that $\Delta\in[0,\frac{1}{4}]$. By taking $\Delta=\min\left(\frac{1}{4},\sqrt{\frac{C_{\infty}^{\star}}{20n}}\right)$ yields concludes Theorem 2. ∎
Appendix C Proof for the Application of Offline Learning for Cascading Bandits
@@
For the monotonicity condition (1), the 1-norm TPM condition (2), Lemma 1 in Wang and Chen [9] yields $B_{1}=1$.
@@
Appendix D Algorithm and Proof for the LLM Cache Application
D.1 Offline Learning for the LLM Cache under the Standard CMAB-T View
@@
if the experimenter samples empty cache in each round as in [34] so that $\nu(q)=1$, it holds that $C_{1}^{*}\leq 2m$ and
@@
D.2 Improved Offline Learning for the LLM Cache by Leveraging the Full-feedback Property and the Vector-valued Concentration Inequality
@@
if the experimenter samples empty cache in each round as in [34] so that $\nu(q)=1$, it holds that $C_{1}^{*}\leq m$ and
@@
D.3 Online learning for LLM Cache
@@
Similar to Zhu et al. [34], we consider the streaming setting where the cache of size $k$ is the only space we can save the query’s response. That is, after we receive query $q_{t}$ each round, if the cache misses the current cache $\mathcal{M}_{t}$, then we can choose to update the cache $\mathcal{M}_{t}$ by adding the current query and response to the cache, and replacing the one of the existing cached items if the cache $\mathcal{M}_{t}$ is full. This means that the feasible set ${\mathcal{Q}}_{t+1}$ needs to be a subset of the $\mathcal{M}_{t}\cup q_{t}$, for $t\in[T]$.
@@
The key difference from the traditional CUCB algorithm, where any super arm $S\in\mathcal{S}$ can be selected, is that the feasible future cache $\mathcal{M}_{t+1}$ in round $t+1$ is restricted to $\mathcal{M}_{t+1}\subseteq\mathcal{M}_{t}\bigcup q_{t}$, where $\mathcal{M}_{t}$ is the current cache and the query that comes to the system. This means that we cannot directly utilize the top-$k$ oracle as in line 7 of Algorithm 3 and other online CMAB-T works [9, 28] due to the restricted feasible action set. To tackle this challenge, we design a new streaming procedure (lines 16-22), which leverages the previous cache $\mathcal{M}_{t}$ and newly coming $q_{t}$ to get the top-$k$ queries regarding $\hat{p}_{t}(q)\underaccent{\bar}{c}_{t}(q)$.
@@
Appendix E Proof for the Influence Maximization Application under the Node-level feedback
Proof for LABEL:thm:IM.
@@ Appendix E Proof for the Influence Maximization Application under the Node-level feedback
where inequality (a) is due to the same derivation of Eq. 18, inequality (b) is due to Eq. 153, inequality (c) is due to $B_{1}\leq V$ by Lemma 2 of [9]. For the last inequality (d), since $\sigma(S^{*};G)=\sum_{u\in\mathcal{V}}p_{u}^{*}$ where $p_{u}^{*}$ is the probability node $u$ is triggered by the optimal action $S^{*}$ and $p^{\mathbb{D}_{\text{arm},S^{*}}}_{u,v}=p^{*}_{u}$, we have $\sum_{(u,v)\in\mathcal{E}}p_{uv}^{\mathbb{D}_{\text{arm}},S^{*}}\leq d_{\max}\sum_{u\in\mathcal{V}}p^{*}_{u}=d_{\max}\sigma(S^{*};G)$, where $d_{\max}$ is the maximum out-degree.
@@
Appendix F Auxiliary Lemmas
@@
The following lemma is extracted from Theorem 14.2 in Lattimore and Szepesvári [123].
@@
For $\Pr\{\mathcal{E}_{arm,1}(u)\}\geq 1-2\delta$, $\Pr\{\mathcal{E}_{arm,2}(\bar{v})\}\geq 1-2\delta$ they are extracted from Lemma 8 from Liu et al. [29] without taking union bound on $t,u,v$ as in Liu et al. [29]. For $\Pr\{\mathcal{E}_{arm,3}(\bar{u},\bar{v})\}\geq 1-2n\delta$, it is extracted from Lemma 8 from Liu et al. [29] by only taking union bound on $n_{0,\bar{u}}$. For $\Pr\{\mathcal{E}_{counter}(u)\}\geq 1-\delta$, $\Pr\{\mathcal{E}_{emp,1}(\bar{v})\}\geq 1-\delta$, $\Pr\{\mathcal{E}_{emp,2}(\bar{u},\bar{v})\}\geq 1-\delta$, they follow the proof of Lemma 6 without taking union bound on $i\in[m]$. ∎
Appendix G Detailed Experiments
@@
G.1 Offline Learning for Cascading Bandits
@@
We evaluate our algorithm (Algorithm 5) in the cascading bandit scenario by comparing it against the following baseline methods: 1. CUCB-Offline [8], the offline variant of the non-parametric CUCB algorithm, adapted for our setting. We refer to this modified version as CUCB-Offline. 2. EMP [119], which always selects the action based on the empirical mean of rewards.
@@
Synthetic Dataset. We conduct experiments on cascading bandits for the online learning-to-rank application described in Section 4.1, where the objective is to select $K=5$ items from a set of $m=100$ to maximize the reward. To simulate the unknown parameter $\mu_{i}$, we draw samples from a uniform distribution $U[0,1]$ over the interval $[0,1]$. In each round $t$ of the offline pre-collected dataset, a ranked list $S_{t}=(a_{t,1},\ldots,a_{t,K})\subseteq[m]$ is randomly selected. The outcome $X_{t,i}$ for each $i\in S_{t}$ is generated from a Bernoulli distribution with mean $\mu_{i}$. The reward at round $t$ is set to 1 if there exists an item $a_{t,k}$ with index $k$ such that $X_{t,a_{t,k}}=1$. In this case, the learner observes the outcomes for the first $k$ items of $S_{t}$. Otherwise, if no such item exists, the reward is 0, and the learner observes all $K$ item outcomes as $X_{t,i}=0$ for $i\in S_{t}$. Fig. 1(a) presents the average suboptimality gaps of the algorithms across different ranked lists $n$. The proposed CLCB algorithm outperforms the baseline methods, achieving average reductions in suboptimality gaps of 47.75% and 20.02%, compared to CUCB-Offline and EMP algorithms, respectively. These results demonstrate the superior performance of CLCB in offline environments.
@@
Real-World Dataset. We conduct experiments on a real-world recommendation dataset, the Yelp datasethttps://www.yelp.com/dataset, which is collected by Yelp [124]. On this platform, users contribute reviews and ratings for various businesses such as restaurants and shops. Our offline data collection process is as follows: we select a user and randomly draw 200 items (e.g., restaurants or shops) that the user has rated as candidates for recommendation. The agent (i.e., the recommender system) attempts to recommend at most $K$ items to the user to maximize the probability that the user is attracted to at least one item in the recommended list. Each item has an unknown probability $\mu_{i}$, derived from the Yelp dataset, indicating whether the user finds it attractive. Regarding feedback, the agent collects cascading user feedback offline, observing a subset of the chosen $K$ items until the first one is marked as attractive (feedback of 1). If the user finds none of the items in the recommended list $S_{t}$ attractive, the feedback is 0 for all items. Fig. 1(b) shows the average suboptimality gaps of different algorithms over $n=100$ rounds across two action sizes ($K=4,8$). Notably, as $K$ changes, the optimal reward also adjusts according to the expected reward $r(S_{t};\bm{\mu})=1-\prod_{i\in S_{t}}(1-\mu_{i}),$ which explains why the suboptimality gap for smaller $K$ tends to be larger compared to that for larger $K$. CLCB achieves the lowest suboptimality gap compared to CUCB and EMP algorithms, demonstrating its strong performance even on real-world data.
G.2 Offline Learning for LLM Cache
@@
In the LLM Cache scenario, we compare our algorithm (Algorithm 3) against two additional baselines: LFU (Least Frequently Used) [34] and Least Expected Cost (LEC) [34].
@@
Synthetic Dataset. For the LLM cache application described in Section 4.2, we simulate the scenario using 100 distinct queries and set the cache size to 40. Consistent with [34], the frequency distribution follows a power law with $\alpha=0.9$, and the ground truth cost for each query processed is drawn from a Bernoulli distribution with parameter 0.5. The simulation is repeated 20 times to ensure robustness, and we report the mean and standard deviation of the results across different dataset sizes $n=\{2,4,8,16,32,64,128,256,512,1024,2048\}$ in Fig. 3(a). Our normalized results suggest that CLCB-LLM-C significantly outperforms the baseline algorithms, LFU and LEC, achieving an average improvement of 1.32$\times$. These results highlight the effectiveness of CLCB-LLM-C in optimizing cache performance for LLM applications.
@@
Real-World Dataset. We use the SciQ dataset [120], which covers a variety of topics, including physics, chemistry, and biology, to evaluate the performance of our proposed CLCB-LLM-C algorithm using OpenAI’s LLMs. The cost is defined as the price for API calls, based on OpenAI’s official API pricing. Since the cost heavily depends on the token count of the input text, we utilize OpenAI’s tiktoken library, designed to tokenize text for various GPT models. We consider two different LLMs with distinct encoding strategies. Specifically, we use GPT-4-o with the “o200k_base” encoding to present the main experimental results. Additionally, we experiment with another variant, GPT-4-turbo, which employs the “cl100k_base” encoding [121]. For the evaluation, we work with 100 distinct prompts from the SciQ dataset in an offline setting, performing a total of 10,000 queries with cache sizes of $K=10$ and $K=20$, respectively. Fig. 3(b) presents the normalized suboptimality gap of cost over $n=100$ rounds. CLCB-LLM-C achieves 36.01% and 20.70%, less cost compared to LFU and LEC, respectively. Moreover, a larger $K$ shows a lower suboptimality gap, which is consistent with Theorem 3. In addition to the results presented in the main text using GPT-4-o with the “o200k_base” encoding, we experiment with another LLM, GPT-4-turbo with the “cl100k_base” encoding. Fig. 4 demonstrates the robustness of our algorithm across different LLMs.

### v2

@@ Appendix D Proof for the Upper Bound Result
Appendix E Proof for the Lower Bound Result
@@
Recall that $A(\mathcal{D})$ is the action returned by algorithm $A$ and we use the Le Cam’s method (Le Cam, 2012):
@@
where inequality (a) is due to Lemma 8 and inequality (b) comes from: $\mathrm{KL}\left(\mathbb{P}_{\bm{p}\otimes\bm{\mu}_{1}}\|\mathbb{P}_{\bm{p}\otimes\bm{\mu}_{2}}\right)\leq\frac{n\mathrm{KL}\left(\mathbb{P}_{\bm{\mu}_{1}}\|\mathbb{P}_{\bm{\mu}_{2}}\right)}{C_{\infty}^{\star}}\leq\frac{n(2\Delta)^{2}}{C_{\infty}^{\star}\left(1/4-\Delta^{2}\right)}\leq 20n\Delta^{2}/C_{\infty}^{\star}$. Here we use the fact that each arm in the path are fully dependent Bernoulli random variables, $\mathrm{KL}\left(\mathrm{Bern}(p)||\mathrm{Bern}(q)\right)\leq\frac{(p-q)^{2}}{q(1-q)}$ and that $\Delta\in[0,\frac{1}{4}]$. By taking $\Delta=\min\left(\frac{1}{4},\sqrt{\frac{C_{\infty}^{\star}}{20n}}\right)$ concludes Theorem 2. ∎
Appendix F Proof for the Application of Offline Learning for Cascading Bandits
@@
For the monotonicity condition (1), the 1-norm TPM condition (2), Lemma 1 in Wang & Chen (2017) yields $B_{1}=1$.
@@
Appendix G Algorithm and Proof for the LLM Cache Application
G.1 Offline Learning for the LLM Cache under the Standard CMAB-T View
2:  for arm $q\in{\mathcal{Q}}$ do
6:  end for
8:  Return: $\hat{\mathcal{M}}$.
@@ G.1 Offline Learning for the LLM Cache under the Standard CMAB-T View
if the experimenter samples empty cache in each round as in (Zhu et al., 2023) so that $\nu(q)=1$, it holds that $C_{1}^{*}\leq 2m$ and
@@
G.2 Improved Offline Learning for the LLM Cache by Leveraging the Full-feedback Property and the Vector-valued Concentration Inequality
@@
if the experimenter samples empty cache in each round as in (Zhu et al., 2023) so that $\nu(q)=1$, it holds that $C_{1}^{*}\leq m$ and
@@
G.3 Online learning for LLM Cache
@@
Similar to Zhu et al. (2023), we consider the streaming setting where the cache of size $k$ is the only space we can save the query’s response. That is, after we receive query $q_{t}$ each round, if the cache misses the current cache $\mathcal{M}_{t}$, then we can choose to update the cache $\mathcal{M}_{t}$ by adding the current query and response to the cache, and replacing the one of the existing cached items if the cache $\mathcal{M}_{t}$ is full. This means that the feasible set ${\mathcal{Q}}_{t+1}$ needs to be a subset of the $\mathcal{M}_{t}\cup q_{t}$, for $t\in[T]$.
@@
The key difference from the traditional CUCB algorithm, where any super arm $S\in\mathcal{S}$ can be selected, is that the feasible future cache $\mathcal{M}_{t+1}$ in round $t+1$ is restricted to $\mathcal{M}_{t+1}\subseteq\mathcal{M}_{t}\bigcup q_{t}$, where $\mathcal{M}_{t}$ is the current cache and the query that comes to the system. This means that we cannot directly utilize the top-$k$ oracle as in line 7 of Algorithm 2 and other online CMAB-T works (Wang & Chen, 2017; Liu et al., 2023b) due to the restricted feasible action set. To tackle this challenge, we design a new streaming procedure (lines 16-22), which leverages the previous cache $\mathcal{M}_{t}$ and newly coming $q_{t}$ to get the top-$k$ queries regarding $\hat{p}_{t}(q)\underaccent{\bar}{c}_{t}(q)$.
@@
Appendix H Proof for the Influence Maximization Application under the Node-level feedback
Proof for Theorem 4.
@@ Appendix H Proof for the Influence Maximization Application under the Node-level feedback
where inequality (a) is due to the same derivation of Eq. 20, inequality (b) is due to Eq. 155, inequality (c) is due to $B_{1}\leq V$ by Lemma 2 of (Wang & Chen, 2017). For the last inequality (d), since $\sigma(S^{*};G)=\sum_{u\in\mathcal{V}}p_{u}^{*}$ where $p_{u}^{*}$ is the probability node $u$ is triggered by the optimal action $S^{*}$ and $p^{\mathbb{D}_{\text{arm},S^{*}}}_{u,v}=p^{*}_{u}$, we have $\sum_{(u,v)\in\mathcal{E}}p_{uv}^{\mathbb{D}_{\text{arm}},S^{*}}\leq d_{\max}\sum_{u\in\mathcal{V}}p^{*}_{u}=d_{\max}\sigma(S^{*};G)$, where $d_{\max}$ is the maximum out-degree.
@@
Appendix I Auxiliary Lemmas
@@
The following lemma is extracted from Theorem 14.2 in Lattimore & Szepesvári (2020).
@@
For $\Pr\{\mathcal{E}_{arm,1}(u)\}\geq 1-2\delta$, $\Pr\{\mathcal{E}_{arm,2}(\bar{v})\}\geq 1-2\delta$ they are extracted from Lemma 8 from Liu et al. (2022) without taking union bound on $t,u,v$ as in Liu et al. (2022). For $\Pr\{\mathcal{E}_{arm,3}(\bar{u},\bar{v})\}\geq 1-2n\delta$, it is extracted from Lemma 8 from Liu et al. (2022) by only taking union bound on $n_{0,\bar{u}}$. For $\Pr\{\mathcal{E}_{counter}(u)\}\geq 1-\delta$, $\Pr\{\mathcal{E}_{emp,1}(\bar{v})\}\geq 1-\delta$, $\Pr\{\mathcal{E}_{emp,2}(\bar{u},\bar{v})\}\geq 1-\delta$, they follow the proof of Lemma 6 without taking union bound on $i\in[m]$. ∎
Appendix J Detailed Experiments
@@
J.1 Offline Learning for Cascading Bandits
@@
We evaluate our algorithm (Algorithm 4) in the cascading bandit scenario by comparing it against the following baseline methods: 1. CUCB-Offline (Chen et al., 2016), an offline variant of the non-parametric CUCB algorithm, adapted for our setting. We refer to this modified version as CUCB-Offline. 2. EMP (Liu et al., 2021), which always selects the action based on the empirical mean of rewards.
@@
Synthetic Dataset. We conduct experiments on cascading bandits for the online learning-to-rank application described in Section 4.1, where the objective is to select $K=5$ items from a set of $m=100$ to maximize the reward (Dai et al., 2025a). To simulate the unknown parameter $\mu_{i}$, we draw samples from a uniform distribution $U[0,1]$ over the interval $[0,1]$. In each round $t$ of the offline pre-collected dataset, a ranked list $S_{t}=(a_{t,1},\ldots,a_{t,K})\subseteq[m]$ is randomly selected. The outcome $X_{t,i}$ for each $i\in S_{t}$ is generated from a Bernoulli distribution with mean $\mu_{i}$. The reward at round $t$ is set to 1 if there exists an item $a_{t,k}$ with index $k$ such that $X_{t,a_{t,k}}=1$. In this case, the learner observes the outcomes for the first $k$ items of $S_{t}$. Otherwise, if no such item exists, the reward is 0, and the learner observes all $K$ item outcomes as $X_{t,i}=0$ for $i\in S_{t}$. Fig. 1(a) presents the average suboptimality gaps of the algorithms across different ranked lists $n$. The proposed CLCB algorithm outperforms the baseline methods, achieving average reductions in suboptimality gaps of 47.75% and 20.02%, compared to CUCB-Offline and EMP algorithms, respectively. These results demonstrate the superior performance of CLCB in offline environments.
@@
Real-World Dataset. We conduct experiments on a real-world recommendation dataset, the Yelp dataset [^4] https://www.yelp.com/dataset, which is collected by Yelp (Dai et al., 2024b). On this platform, users contribute reviews and ratings for various businesses such as restaurants and shops. Our offline data collection process is as follows: we select a user and randomly draw 200 items (e.g., restaurants or shops) that the user has rated as candidates for recommendation. The agent (i.e., the recommender system) attempts to recommend at most $K$ items to the user to maximize the probability that the user is attracted to at least one item in the recommended list. Each item has an unknown probability $\mu_{i}$, derived from the Yelp dataset, indicating whether the user finds it attractive. Regarding feedback, the agent collects cascading user feedback offline, observing a subset of the chosen $K$ items until the first one is marked as attractive (feedback of 1). If the user finds none of the items in the recommended list $S_{t}$ attractive, the feedback is 0 for all items. Fig. 1(b) shows the average suboptimality gaps of different algorithms over $n=100$ rounds across two action sizes ($K=4,8$). Notably, as $K$ changes, the optimal reward also adjusts according to the expected reward $r(S_{t};\bm{\mu})=1-\prod_{i\in S_{t}}(1-\mu_{i}),$ which explains why the suboptimality gap for smaller $K$ tends to be larger compared to that for larger $K$. CLCB achieves the lowest suboptimality gap compared to CUCB and EMP algorithms, demonstrating its strong performance even on real-world data.
J.2 Offline Learning for LLM Cache
@@
In the LLM Cache scenario, we compare our algorithm (Algorithm 2) against two additional baselines: LFU (Least Frequently Used), which is a caching strategy that evicts the least frequently accessed items to optimize cache usage (Zhu et al., 2023), and Least Expected Cost (LEC), which is an advanced caching algorithm that minimizes inference cost by evicting items with the lowest estimated expected cost (Zhu et al., 2023).
@@
Synthetic Dataset. For the LLM cache application described in Section 4.2, we simulate the scenario using 100 distinct queries and set the cache size to 40. Consistent with (Zhu et al., 2023), the frequency distribution follows a power law with $\alpha=0.9$, and the ground truth cost for each query processed is drawn from a Bernoulli distribution with parameter 0.5. The simulation is repeated 20 times to ensure robustness, and we report the mean and standard deviation of the results across different dataset sizes $n=\{2,4,8,16,32,64,128,256,512,1024,2048\}$ in Fig. 3(a). Our normalized results suggest that CLCB-LLM-C significantly outperforms the baseline algorithms, LFU and LEC, achieving an average improvement of 1.32$\times$. These results highlight the effectiveness of CLCB-LLM-C in optimizing cache performance for LLM applications.
@@
Real-World Dataset. We use the SciQ dataset (Welbl et al., 2017), which covers a variety of topics, including physics, chemistry, and biology, to evaluate the performance of our proposed CLCB-LLM-C algorithm using OpenAI’s LLMs. The cost is defined as the price for API calls, based on OpenAI’s official API pricing. Since the cost heavily depends on the token count of the input text, we utilize OpenAI’s tiktoken library, designed to tokenize text for various GPT models. We consider two different LLMs with distinct encoding strategies. Specifically, we use GPT-4-o with the “o200k_base” encoding to present the main experimental results. Additionally, we experiment with another variant, GPT-4-turbo, which employs the “cl100k_base” encoding (OpenAI, 2025). For the evaluation, we work with 100 distinct prompts from the SciQ dataset in an offline setting, performing a total of 10,000 queries with cache sizes of $K=10$ and $K=20$, respectively. Fig. 3(b) presents the normalized suboptimality gap of cost over $n=100$ rounds. CLCB-LLM-C achieves 36.01% and 20.70%, less cost compared to LFU and LEC, respectively. Moreover, a larger $K$ shows a lower suboptimality gap, which is consistent with Theorem 3. In addition to the results presented in the main text using GPT-4-o with the “o200k_base” encoding, we experiment with another LLM, GPT-4-turbo with the “cl100k_base” encoding. Fig. 4 demonstrates the robustness of our algorithm across different LLMs.